OpenStudy (anonymous):

2nd order ODE problem.

OpenStudy (anonymous):

I am doing something stupidly wrong here and I see where it is

OpenStudy (anonymous):

I got the characteristic question to be \[r ^{2}-r+4=0\]

OpenStudy (anonymous):

No wait r^2-4r+4=0

OpenStudy (anonymous):

Nevermind I just realised I am looking at a trinomial. roots will be the same I believe.

OpenStudy (anonymous):

you found the solutions yet? @ironictoaster ?

OpenStudy (anonymous):

there is only one from what I can tell r=2

OpenStudy (anonymous):

Yeah I got that, general solution seems to be y=c1*e^2x+c2*e^2x

OpenStudy (anonymous):

Derive that equation and use my initial conditions.

OpenStudy (anonymous):

almost from what I can see here, you have to add an additional x to one of them, that's how it works for multiply identical roots. \[y=\large c_1e^{2x}+c_2xe^{2x} \]

OpenStudy (anonymous):

I did not know that. I've never done question that had double roots.

OpenStudy (anonymous):

So that derived would be y'=2c1e^2x+2c1xe^2x

OpenStudy (anonymous):

I really should learn latex..

OpenStudy (anonymous):

I would first set f(0)=1 that leads to c_1 =1 right?

OpenStudy (anonymous):

so you can forget about that constant there.

OpenStudy (anonymous):

Sorry the second constant should be c\[_{2}\]

OpenStudy (anonymous):

\[c_{2}\]

OpenStudy (anonymous):

I get the following as a derivative \[\Large y'=c_12e^{2x}+ c_2e^{2x}+c_22xe^{2x} \]

OpenStudy (anonymous):

Yeah I got the same, I just pulled out the 2's in front of the terms.

OpenStudy (anonymous):

got, so you can now substitute your initial conditions and you are done, from y(0) follows directly that c_1 = 1 from what I see, so I would use that to your advantage

OpenStudy (anonymous):

Yeah I am left with two similtanous equations. to solve for c1 and c2

OpenStudy (anonymous):

\[c_{1}+c _{2}=1\] \[2c _{1}+c _{2}=4\]

OpenStudy (anonymous):

2c1+2c2=4***

OpenStudy (anonymous):

I don't think this can be done, everything just cancels out....

OpenStudy (anonymous):

the first one doesn't seem correct to me, if you substitute y(0) you are left with c_1=1

OpenStudy (anonymous):

\[ y=\large c_1e^{2x}+c_2xe^{2x}\ \\ y(0)=c_1=1\]

OpenStudy (anonymous):

second term includes an additional x, therefore it falls away

OpenStudy (anonymous):

Awh I forgot to add in the x when writing the equation down. Yeah multiplying by 0 cancels the terms out.

OpenStudy (anonymous):

the 2nd term that is

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yeah the missing x was screwing me up, thanks for the help, the little mistakes will be the end of me!

OpenStudy (anonymous):

hehe okay great!